### 08 Mar 2024 v3

# https://youjiazhang.github.io/USD/ Optimized View and Geometry Distillation from Multi-view Diffuser

Abstract
Generating multi-view images from a single input view using image-conditioned diffusion models is a recent advancement and has shown considerable potential. However, issues such as the lack of consistency in synthesized views and over-smoothing in extracted geometry persist. Previous methods integrate multi-view consistency modules or impose additional supervisory to enhance view consistency while compromising on the flexibility of camera positioning and limiting the versatility of view synthesis. In this study, we consider the radiance field optimized during geometry extraction as a more rigid consistency prior, compared to volume and ray aggregation used in previous works. We further identify and rectify a critical bias in the traditional radiance field optimization process through score distillation from a multi-view diffuser. We introduce an Unbiased Score Distillation (USD) that utilizes unconditioned noises from a 2D diffusion model, greatly refining the radiance field fidelity. We leverage the rendered views from the optimized radiance field as the basis and develop a two-step specialization process of a 2D diffusion model, which is adept at conducting object-specific denoising and generating high-quality multi-view images. Finally, we recover faithful geometry and texture directly from the refined multi-view images. Empirical evaluations demonstrate that our optimized geometry and view distillation technique generates comparable results to the state-of-the-art models trained on extensive datasets, all while maintaining freedom in camera positioning.
1 Introduction
Traditionally, the process of generating a three-dimensional model from a singular image necessitates extensive and meticulous efforts by highly skilled artists. However, recent advancements in neural networks, particularly through the adaptation of 2D diffusion models for 3D synthesis, have rendered the conversion of a single image into a 3D object feasible. The early breakthrough comes from the text to 3D domain, where DreamFusion [30] and Score Jacobian Chaining (SJC) [46] proposes a Score Distilling Sampling (SDS) strategy to distill the scores learned by 2D diffusion models from large-scale images to optimize a Neural Radiance Field (NeRF) [27], circumventing the need for 3D data. Successive approaches further improve the quality and diversity of generated geometries from textural prompts [48, 19, 2]. Particularly, RealFusion [26] migrates the scheme to generate plausible 3D reconstruction matches to a single input image via textual inversion adapted supervision.
More relevantly, 3DiM [50] and MVDream [38] develop a pose-conditional image-to-image diffusion model, which generates the novel view at a target pose from a source view. Zero-1-to-3 [21] adopts a similar framework and learns control of viewpoints through a synthetic dataset and demonstrates zero-shot generalization to in-the-wild images. Though Zero-1-to-3 demonstrates plausible novel views, they are not multi-view consistent and the geometry distilled from SDS tends to be oversmoothed. To enhance the multi-view consistency, SyncDreamer [23] devises a volume-encoded multi-view noise predictor to share information across different views. Wonder3D [24] predicts the multi-view color images along with their normal maps from a cross-domain diffusion model. Though enhancing the multi-view consistency of image generation, these methods compromise the flexibility of camera positioning and only allow synthesis for a limited number of views.
In this study, we observe that the predicted unconditional noise from the multi-view diffuser, i.e., the Zero-1-to-3 model, appears to be biased. That is, even if we only add very low-level noise to a normal image and use the unconditional noise predicted by a Zero-1-to-3 model for denoising, the result still tends to deviate greatly from the original image. We analyze and rectify the critical bias by using an unconditioned noise from a pre-trained 2D diffusion model and greatly refining the geometry fidelity. Moreover, previous approaches use either 3D volume [23] or ray aggregation [45] to share information across views. We consider the radiance field as the consistency prior, and encourage the generated multi-view images to be consistent with the NeRF renderings. We develop a two-step specialization process of a 2D diffusion model, which is adept at conducting target-specific denoising and generating high-quality multi-view images from NeRF renderings. We then further use the refined views to generate the geometry and texture from NeuS [47], and in the meanwhile enforce input view consistency using view score distillation. Our approach generates comparable-quality of multi-view images and geometry to the SOTA approaches, including SyncDreamer and Wonder3D, without enforcing any restriction on camera poses. Consequently, we posit that our approach offers superior adaptability and effectiveness in addressing the challenges associated with generating consistent and high-quality multi-view imagery and geometry, affirming its substantial potential for widespread application in relevant fields.
2 Related Work
2D Diffusion Models for 3D Generation. Recent advancements in 2D diffusion models [13, 34, 4] and large-scale visual language models, notably the CLIP model [33], have catalyzed new approaches for generating 3D assets. Pioneering efforts such as DreamFusion [30] and SJC [46] have developed methods for transforming 2D text into images, subsequently facilitating the generation of 3D shapes from text. This approach has inspired a range of subsequent studies that adopt a shape-by-shape optimization scheme. Additionally, the integration of 2D diffusion models with robust vision language models, especially CLIP [33], has emerged as a significant exploration in the generation of 3D assets [54]. The typical methodology involves optimizing a 3D representation, such as NeRF, mesh, or SDF, and then utilizing neural rendering to generate 2D images from various viewpoints. These images are processed through 2D diffusion models or the CLIP model to calculate SDS losses, which guide the optimization of the 3D shape. Building on the foundations laid by DreamFusion and SJC, numerous works have enhanced text-to-3D distillation methods in various aspects. Notably, Magic3D [19] develops a two-stage coarse-to-fine optimization framework for high-resolution 3D content generation, and ProlificDreamer [48] proposes a Variational Score Distillation (VSD) for generating highly detailed geometry. However, challenges such as low efficiency and the multi-face Janus problem, where optimized geometry tends to produce multiple faces due to the lack of explicit 3D supervision, remain prevalent. Furthermore, some works [32] have applied this distillation pipeline in single-view reconstruction tasks. While these methods have achieved impressive results, they often require extensive time for textual inversion and NeRF optimization, without always guaranteeing satisfactory outcomes. In contrast with the 2D diffusion to 3D extension, which is ignorant to multi-view consistency, our method focuses on the multi-view diffuser technique, which predicts the noises for novel views, and inherently avoids the Janus problem.
Multi-view Diffusion Models. In light of the complexities involved in ensuring the integrity of generated 3D content, recent efforts [50, 9, 6, 60, 45, 1, 57, 43, 56, 40, 42, 52, 22, 18] have explored the feasibility of directly generating novel views from a single image input. Notably, the Scene Representation Transformer [37] extends the vision transformer to image sets, enabling global information integration for 3D reasoning. Similarly, 3DiM [50] develops a pose-conditional image-to-image diffusion model, translating a single input view into consistent and sharp completions across multiple views. A seminal work in this area, Zero-1-to-3 [21], utilizes a similar network structure, trained on a large-scale synthetic 3D dataset, demonstrating notable generalizability. Recent advancements, such as One-2-3-45 [53], have leveraged the generalizable neural reconstruction method SparseNeuS [25] to directly produce 3D geometry from images generated by Zero-1-to-3. While this approach is more efficient and alleviates the Janus (multi-head) problem, it tends to produce lower-quality results with reduced geometric detail. In a different vein, SyncDreamer [23] concentrates on object reconstruction, generating images in a single reverse process and utilizing attention to synchronize states among views. This contrasts with Viewset Diffusion [39], which requires predicting a radiance field. SyncDreamer solely relies on attention for synchronization, fixing the viewpoints of generated views to enhance training convergence. A significant trend in recent research has been the design of various functional attention layers. Consistent123 [51] employs a shared self-attention layer, where all views query the same key and value from the input view. ConsistNet [55] introduces two sub-modules: a view aggregation module and a ray aggregation module, to extract features consistent across multiple views. MVDream [38] utilizes 3D self-attention. Wonder3D [24] goes a step further by not only outputting multi-view images but also inputting the normal map for each perspective, using cross-domain attention to maintain the consistency of the 3D structure. Both SyncDreamer and Wonder3D, resulting from fixed view outputs, exhibit sensitivity to the camera angle of the input image. In contrast, our method can reconstruct accurate 3D structures from inputs captured at varying camera angles. While existing approaches have focused on developing novel modules for enforcing multi-view consistency, our work demonstrates that an optimized distillation strategy can yield views and geometries comparable to models trained on large-scale datasets. This insight may inspire further exploration in improved strategies for geometry and view extraction.
3 Method
Our objective is to synthesize consistent multi-view images and high-quality geometric representations from a single input image. Notable prior work, such as the Zero-1-to-3, has demonstrated impressive results by utilizing an image and camera pose-conditioned diffusion model. However, this approach encounters limitations, particularly in terms of generating inconsistent multi-view images and a tendency for over-smoothing in the geometric output. The state-of-the-art research, including SyncDreamer [23] and Wonder3D [24], addresses these challenges by incorporating additional modules for consistency or employing normal supervision. However, this often compromises the flexibility of positioning the target camera at will. Contrastingly, our work adopts a distinct methodology, showcasing that comparable quality in views and geometry can be achieved through a meticulously crafted distillation strategy. Central to our approach is the insight that the unconditional noise predictions from Zero-1-to-3 are inherently biased. We propose the utilization of unconditional noise from the Stable Diffusion [41] model to rectify this issue, as elaborated in Sec. 3.1. Our method, termed Unbiased Score Distillation (USD), significantly enhances the quality of the radiance field relative to previously used SDS/SJC methods. Furthermore, we employ the optimized NeRF [27] as a consistency prior, in contrast to previous implicit constraints such as 3D volume or ray aggregation. We ensure that the generated views and geometry align coherently with the distilled NeRF to achieve consistency. We posit that specializing a diffusion model to denoise the target object is crucial. To this end, we implement the DreamBooth [35] technique and engage in a two-stage fine-tuning process, detailed in Sec. 3.2. The first stage involves using multi-view images from Zero-1-to-3 as positive samples, contrasting them with text-prompt generated images as negative samples for image style learning. In the subsequent stage, the input image serves as the positive sample, with all Zero-1-to-3 generated images as negatives, focusing on learning finer details. Subsequently, we introduce a low level of noise into the NeRF renderings and employ the fine-tuned diffusion model for denoising. The final step involves applying the NeuS technique for mesh reconstruction, while in the meanwhile using View Score Distillation to ensure input view consistency. Fig. 3 illustrates the complete pipeline of our methodology.
3.1 Unbiased Score Distillation
In this section, we highlight the significant bias in Zero-1-to-3’s unconditional noise due to insufficient unconditional training and object-level dataset bias, affecting geometry quality in SDS-based distillation, both theoretically and empirically. We then propose a rectification method.
Bias in Unconditional Noise. As a multi-view diffuser, Zero-1-to-3 predicts noises of a target image latent given two conditions: the input image and relative camera pose , i.e., it learns the probability distribution for image latent . We apply Bayes’ Rule to decompose the conditional probability:
Diffusion models estimate the score [15] of the data distribution, i.e., the derivative of the log probability, giving us the following expression:
This leads to the score estimate with classifier-free guidance (CFG [14]):
where is a neural network for predicting noises given conditions, and are guidance scales that enable separately trading off the strength of conditions and separately. To ensure accurate prediction of and , one needs to randomly drop and during training . However, we observed that Zero-1-to-3 did not truly follow this procedure to supervise unconditional noise. As shown in Fig. 5, Zero-1-to-3 only randomly dropped image conditions while keeping untouched. During inference, it replaces the tensor after fully connected (FC) layer with zeros (not exactly equal to setting both and to zeros as ), leads to a bias. Moreover, the dataset used for fine-tuning is majorly object-centric, which may also introduce additional domain bias. Fully addressing this bias problem requires re-training the multi-view diffuser on a broader and more balanced dataset, which requires tremendous effort. Here we propose an effortless fix in the following.
Rectification. To alleviate the inaccurate unconditional noise problem, we first set to eliminate . We then and yield a special case of as:
To demonstrate our setting of and is valid, We provide an empirical validation as in Fig. 7. Further, we consider predicts noises from only the noisy latent and is equivalent to the unconditional noise of Stable Diffusion (SD) as it uses the same variational autoencoders (VAE [17]) as Zero-1-to-3. As such, the bias has been well rectified. To verify the effect of our rectification, Table 1 shows the denoising effect using various unconditional noise settings: (1) use SD unconditional noise . (2) use Zero-1-to-3 noise with as a condition (same to the training process as in Fig 5). (3) use Zero-1-to-3 noise with as a condition (referring to the inference process, as shown in Fig 5). The SD noise generates the best result, and setting (2) is slightly better than (3) as it follows the training set up.
| Noise level | T=50 | T=100 | T=200 | T=300 | ||||||||
| Metrics | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS | PSNR | SSIM | LPIPS |
| 40.26 | 0.992 | 0.008 | 36.94 | 0.988 | 0.014 | 33.99 | 0.981 | 0.023 | 31.98 | 0.974 | 0.027 | |
| 36.17 | 0.981 | 0.015 | 34.85 | 0.979 | 0.024 | 33.07 | 0.975 | 0.031 | 30.44 | 0.969 | 0.046 | |
| 36.13 | 0.979 | 0.015 | 34.70 | 0.970 | 0.026 | 32.76 | 0.973 | 0.033 | 29.94 | 0.963 | 0.055 | |
Unbiased Score Distillation. One major application of a multi-view diffuser is to distill 3D content, represented as NeRF paramater , using the SDS loss for optimization:
where is a weighting function, is standard Gaussian noise, and refers to the noisy latent as , with being the noise scheduler. We rewrite the noise difference in Formula 5 by adding an additional weighting factor to :
where setting we get Formula 5. Inspired by DDS [12] and CSD [58], we found that setting can significantly improve the details of the 3D details generated using SDS, as shown in Fig. 12. We obtain our Unbiased Score Distillation (USD) as:
Our USD generates much better and consistent 3D than the SDS/SJC method used in Zero-1-to-3, as shown in Fig. 7.
3.2 Consistent View and Geometry Distillation
Though with USD, we can obtain optimized NeRF for view synthesis, the generated images tend to be blurry, and direct extract geometry from the density field of NeRF produces noises on the mesh. Since our ultimate goal is to extract high-quality geometry and consistent multi-views, we propose to utilize the generated NeRF as the view consistency prior, i.e., the final high-quality view images should be consistent with the NeRF rendering. Thus, our problem transformed into a denoising problem.
Two-stage Specified Diffusion. We leverage the recent advance in diffusion model specialization, i.e., the DreamBooth [35], to fine-tune a 2D diffusion model for the specific target in the input image. We observe that the novel view images generated by Zero-1-to-3, though not multi-view consistent, tend to have the same style as the input image. We design a two-stage tuning method to gradually let the diffusion model learn the object details. In the first stage, we use the multi-view images generated by Zero-1-to-3 as positive samples, contrasting them with text-prompt generated images as negative samples for learning the visual style of the target. In the subsequent stage, the input image serves as the positive sample, with all Zero-1-to-3 generated images as negatives, focusing on learning finer details. During optimization, we use a unique identifier [V] to capture the visual style of the target. In the second stage, we set only the input image as the positive sample and the images generated by Zero-1-to-3 as negative ones, and use the additional identifier [S] to capture the identity of the specific target.
Geometry and Texture Distillation with Input View Supervision. With the specialization diffusion model, we add a small noise, Stable Diffusion scheduler , to the NeRF render images and conduct the denoising process. Then, we use the NeuS technique to reconstruct the geometry from the high-quality and clear images.
We observe that the input view is rarely used for optimization of the geometry and texture in multi-view diffusers, despite the fact that the input image is most faithful to the target object. To exploit the input view information, we further develop a reference view score distillation during the NeuS reconstruction process. Specifically, we consider the image render from NeuS at the input image viewpoint as , where is the rendering function from NeuS model defined by , is the camera pose of the input image, can be set to a particular relative pose. We define the following reference view distillation loss as:
where is input view image. We add this reference view distillation loss to original photometric loss with optimized images for optimizing the NeuS parameter . This scheme is a better supervision strategy than directly applying MSE loss on the input view.
4 Experiments
We conduct extensive experiments, both qualitatively and quantitatively, to demonstrate the effectiveness of our method.
4.1 Implementation details
Optimizing the pipeline. We use exactly the same set of hyperparameters for all experiments and do not perform any per-object hyperparameter optimization. We use the implicit-volume implementation in threestudio [11] as our 3D representation(NeRF), which includes a multi-resolution hash-grid and a MLP to predict density and RGB. The NeRF is optimized for 10,000 steps with an Adam optimizer at a learning rate of 0.01, weight decay of 0.05, and betas of (0.9, 0.95). For USD, the maximum and minimum time steps are decreased from 0.98 to 0.5 and 0.02, respectively, over the first 5,000 steps. We adopt the Stable Diffusion [41] model of V2.1. The classifier-free guidance (CFG) scale of the USD is set to 7.5 following [48]. The DreamBooth backbone is implemented by Stable Diffusion V2.1. In the first stage, we use Stable Diffusion to generate 200 images as negative samples for training. The USD process takes about 1.5 hours on a NVIDIA Tesla V100 (32GB) GPU.For DreamBooth fine-tuning, we train the model around for 600 steps with a learning-rate as 2e-6, weight decay as 0.01 and a batch size of 2.
Camera setting. Since the reference image is unposed, we assume its camera parameters [8] are as follows. We set the field of view (FOV) of the camera is 40, and the radial distance is 1.5 meters. Note this camera setting works for images subject to the front-view assumption. For images taken deviating from the front view, a manual change of polar angle or a camera estimation is required.
4.2 Evaluation Protocol
Evaluation Datasets. Following prior research [21, 23, 24], we adopt the Google Scanned Object dataset [7] for our evaluation, which includes a wide variety of common everyday objects. Our evaluation dataset matches that of SyncDreamer [23], consisting of 30 objects that span from everyday items to animals. For each object in the evaluation set, we render an image with a size of 256×256 and use it as the input. Additionally, to assess the generalization ability of our model, we include images with diverse styles collected from the website in Zero-1-to-3, SyncDreamer and Wonder3D.
| Method | Chamfer Dist. | Volume IoU |
| Realfusion [26] | 0.0819 | 0.2741 |
| Magic123 [31] | 0.0516 | 0.4528 |
| One-2-3-45 [20] | 0.0629 | 0.4086 |
| Point-E [29] | 0.0426 | 0.2875 |
| Shap-E [16] | 0.0436 | 0.3584 |
| Zero-1-to-3 [21] | 0.0339 | 0.5035 |
| SyncDreamer [23] (NeuS) | 0.0261 | 0.5421 |
| Wonder3D [24] (iNGP+NeuS) | 0.0199 | 0.6244 |
| Ours (NeuS) | 0.0240 | 0.5688 |
| Ours (iNGP+NeuS) | 0.0177 | 0.6330 |
Metrics. To evaluate the quality of single view reconstruction, we used two commonly used metrics: the chamfer distance (CD) between the ground truth shape and the reconstructed shape, and the volume IoU. Due to different methods using different normative systems, before calculating these two metrics, we first align the generated shapes with the basic fact shapes. Moreover, we adopt the metrics PSNR, SSIM [49] and LPIPS [59] for evaluating the generated color images.
4.3 Ablation Study
We validate our design choices by ablating 4 major model variants, that are without Unbiased Score Distillation, without reference view distillation loss, without the two-stage DreamBooth(DB). The texture color of the 3D models geneDirated by without USD exhibit bias, and the shapes are not smooth. Not using reference view supervision will result in the inability to recover the same texture details as the input image, and not using DB will result in blurring of texture details.
4.4 Different Viewing Angle Comparisons
4.5 Single View Reconstruction
We evaluate the quality of the reconstructed geometry of different methods. The quantitative results are summarized in Table 3, and the qualitative comparisons are presented in Fig. 9. The quality of Wonder3D [24] shape reconstruction depends on the perspective of the input view, such as ‘Train’ shown in Fig. 9, where Wonder3D generated incorrect prediction results. The shape generated by SyncDreamer [23] undergoes deformation due to the camera pose in the input view. One-2-3-45 [20] attempts to reconstruct meshes from the multiview-inconsistent outputs of Zero-1-to-3 [21]. While it can capture coarse geometries, it loses important details in the process. In contrast, our method can achieve good reconstruction quality and texture in terms of geometric structure and texture. In the paper, all the surface extraction demonstrated by our method is built on the Instant NGP(iNGP) [28] based SDF reconstruction method [10], and we use Blender Cycles [3] to render the results.
4.6 Novel View Synthesis
We evaluate the quality of novel view synthesis for different methods. The quantitative results are presented in Table 3, and the qualitative results can be found in Fig. 4. Zero-1-to-3 [21] produces visually reasonable images, but they lack multi-view consistency since it operates on each view independently (we don’t show the results of Zero-1-to-3). Although SyncDreamer [23] introduces a volume attention scheme to enhance the consistency of multi-view images, their model is sensitive to the elevation degrees of the input images and tends to produce unreasonable results. Wonder3D [24] ensures 3D consistency by generating normal maps, but may result in incorrect results for some input viewpoints.
4.7 Text-to-image-to-3D
As a case study, we combine text-to-image models, i.e., the Stable Diffusion or Imagen [36] to generate 3D models from text. We show some examples in Fig. 12. Compared to DreamFusion [30] and ProlificDreamer [48], our method shows no multi-face Janus problem and conforms to the text faithfully.
5 Conclusions
In this work, we introduce an optimized approach for distilling geometry and views from a multi-view diffuser, with a specific focus on the Zero-1-to-3 model. We observed that the direct application of the SDS/SJC technique to Zero-1-to-3 is often suboptimal, primarily due to bias issues inherent in unconditional noise. To address this, we propose an Unbiased Score Distillation (USD) strategy by leveraging unconditioned noises from a 2D diffusion model to effectively enhance the optimized radiance field. Moreover, we developed a two-stage DreamBooth refinement process to improve the rendering of views. This process ensures consistency across multiple perspectives while simultaneously enhancing image quality. While we have identified and addressed the bias issue in the Zero-1-to-3 model, the underlying causes remain to be fully understood. Future research will delve into the theoretical aspects of this bias problem. Additionally, we aim to explore the potential applications of USD in other fields, such as image translation and view synthesis.